# Domain decomposition of the modified Born series approach for large-scale wave propagation simulations

### v1

@@ Abstract
The modified Born series method is a fast and accurate method for simulating wave propagation in complex structures. Currently, its main limitation is that the size of the simulation is limited by the working memory of a single computer or graphics processing unit (GPU). Here, we present a domain decomposition method that removes this limitation. We show how to decompose large problems over subdomains while maintaining the accuracy, memory efficiency, and guaranteed monotonic convergence of the method. We demonstrate our approach by solving the Helmholtz problem for a complex structure of size $315\!\times\!315\!\times\!$ 315 × 315 wavelengths in just 379 seconds on a dual-GPU system.
@@
Helmholtz equation , Born series , Domain decomposition
journal: Journal of Computational Physics
@@ 1 Introduction
Osnabrugge et al. [3] showed that the MBS could be executed on a graphics processing unit (GPU) efficiently, solving Maxwell’s equations in a structure of $28\!\times\!28\!\times\!$ 28 × 28 wavelengths in just 3 seconds, representing a factor of $50$ speedup compared to performing the simulations on a central processing unit (CPU). Recently, Valantinas and Vettenburg [26] solved the Helmholtz equation for a system of $2.1\cdot 7$ cubic wavelengths by mapping the steps of the MBS to layers in an artificial neural network framework, leveraging the extensive software and hardware acceleration available for these applications. However, GPUs have a limited amount of working memory that is expensive and cannot be expanded as flexibly as the main working memory, adding a severe limitation to the simulation size [27, 28].
@@
Here, we present a domain decomposition approach for the MBS for solving the Helmholtz equation. Our approach maintains the key benefits of MBS: low memory use, high accuracy, and guaranteed convergence, while minimising the overhead on memory use and computation time.
We demonstrate the domain decomposition approach on a dual-GPU system through a large 3D simulation of $3.1\cdot 10^{7}$ cubic wavelengths, a size that is 1.93×1.93$\times$1.93 × larger than the maximum that could be accommodated on a single GPU and only took $379$ seconds to complete. An open-source implementation in Python is available on GitHub [33].
@@ 2 Modified Born series
The key step in both the modified [2] and the original [35] Born series formalisms is to split $A=L+V$, with $L\coloneqq c\left[\nabla^{2}+k_{0}^{2}\right]$ corresponding to wave propagation in a homogeneous medium with background wavenumber $k_{0}$, and $V=c\left[{k}^{2}(\boldsymbol{r})-k_{0}^{2}\right]$ the scattering potential.
@@
In contrast to the original Born series, which diverges for large or strongly scattering media [39], the MBS converges monotonically to the solution of Eq. (1) provided that $\alpha\leq 1$, and $ic$ is a real number less than $1/|k^{2}(\boldsymbol{r})-k_{0}^{2}|_{\infty}$ [2], where $|\cdot|_{\infty}$ denotes the maximum value over all $\boldsymbol{r}$. We can simply truncate this iteration when the norm of the residual $\Gamma^{-1}\left(\boldsymbol{y}-A\boldsymbol{x}^{(n)}\right)$ drops below a certain threshold.
@@
Figure 1: Matrix representation of splitting operator $A$ into $L+V$ for a 1D system. (a) Operator $A$ is split into (b) $L$ containing the Laplacian operator, a band along the diagonal, and (c) $V$, a diagonal matrix containing the scattering potential. (d) The operator $L$ with wraparound artefacts resulting from implementing $L$ as a fast convolution (Eq. (5)). Note that the colour scale is logarithmic in both the positive and negative directions from 00.
The splitting $A=L+V$ is visualised in Fig. 1, where the operators are represented as matrices for illustrative purposes. In practice, these operators are not implemented as matrices since the full matrices would be too large to store in memory. The alternative of using sparse matrices would require replacing $L$ by a matrix with only a few diagonals, which is exactly the finite difference approximation that causes numerical dispersion and thus should be avoided. Instead, $L$ can simply be implemented as a convolution with the angular spectrum kernel [40, 41]
@@
where $\mathcal{F}$ and ${\mathcal{F}}^{-1}$ denote the forward and inverse Fourier transform, respectively, and $\boldsymbol{p}$ denotes the Fourier-space coordinate vector. In practice, a fast Fourier transform is used to evaluate this convolution efficiently. As shown in Fig. 1d, this approach does introduce wrapping artefacts. These artefacts can be greatly reduced by using an alternating sequence of offset Fourier transforms [3] and completely eliminated using our proposed method (see Section 3).
@@ 3 Domain decomposition of the modified Born series
An essential generalisation of the MBS introduced in [34] is that we have full freedom to choose the splitting $A=L+V$, as long as the scaling factor $c$ can be chosen such that $\lVert V\rVert<1$. We use this freedom to implement our domain decomposition framework.
@@
This freedom enables us to implement $L$ as a fast convolution over each subdomain (see Fig. 2b), giving in block matrix form
$L=\begin{bmatrix}L_{11}&0\\ 0&L_{22}\end{bmatrix},\quad V=\begin{bmatrix}A_{11}-L_{11}&A_{12}\\ A_{21}&A_{22}-L_{22}\end{bmatrix}.$ (8)
Since by definition $V\coloneqq A-L$, the wraparound artefacts caused by the fast convolution in $L$ appear in $V$ with the opposite sign. As can be seen in Fig. 2c, $V$ now has three contributions: the original scattering potential on the diagonal, the correction for the wrapping artefacts caused by the fast convolution, which we call $C_{11}$ and $C_{22}$, and the off-diagonal blocks $A_{12}$, and $A_{21}$ that represent communication between the subdomains.
Matrices $C_{11}$, $C_{22}$, $A_{12}$, and $A_{21}$ can be constructed directly by comparing two situations: 1) the angular spectrum kernel applied to a point source using a linear convolution (Fig. 3a), and 2) the same kernel applied using a fast convolution on a subdomain (Fig. 3b). The difference between these two curves gives the required wrapping correction (which corresponds to a column of matrix $C_{11}$) and the transfer to the second subdomain (a column of matrix $A_{12}$). By repeating this process for different positions of the point source and thus different columns of the matrices, we recover matrices $C_{11}$ and $A_{12}$. Due to the symmetry of the angular spectrum kernel, the bottom-left part of matrix $C_{11}$ is equal to $-A_{12}$. The top-right part of matrix $C_{11}$ is given by the transpose $-A_{12}^{T}$.
Figure 3c shows the bottom-left part of the domain-transfer matrix $A_{12}$ (also see Fig. 2c). It can be seen that the magnitude of the elements decreases rapidly away from the subdomain border. Therefore, we can truncate matrices $C$ and off-diagonal blocks of $A$ to only take into account $t\ll N$ edge points. This way, the overhead of these additional computations is minimised.
Figure 4: Using truncated corrections in $V$ to reconstruct $A$. (a) Matrix representation of $A$ decomposed over two subdomains, demarcated by the dashed grey lines, but with truncated corrections. (b) $L$ decomposed over the subdomains remains unchanged and is the same as in Fig.2b. (c) $V$ with truncated corrections for wrapping and transfer, visible as square blocks of side $t=6$ in the corners of subdomain blocks.
@@ 3.1 Implementation
The domain decomposition framework is incorporated into the preconditioned Richardson iteration Eq. (6), but with the operators $L$ and $V$ computed for and applied to each subdomain. Thus, the fast convolution with ${(L+I)}^{-1}$ is computed for each subdomain $s$ as ${(L_{s}+I)}^{-1}$. As discussed in Section 3, operator $V$ has three contributions, i.e. the scattering potential on the diagonal, which is now applied over the subdomains, the wrapping corrections $C$, and the transfer corrections (off-diagonal blocks of $A$). The steps for computing and applying the operator $V$ for every subdomain $s$ are as below:
@@
With this activation strategy, for example, a 3D simulation of size $100\!\times\!100\!\times\!$ 100 × 100 wavelengths split into $3\!\times\!1\!\times\!1$ subdomains was $12\%$ faster than without this strategy.
@@ 4 Results and Discussion
In this section, we demonstrate the domain decomposition of the MBS in 3D simulations. For all simulations in this section, we simulate light propagation through a slab of scattering material. The slab extends infinitely in the $y$ and $z$ directions, along which we use periodic boundaries. It is illuminated with a planar source located at $x=0$, and we use absorbing boundaries with a thickness of 5 wavelengths along the $x$ direction [3].
The slab is placed in a medium with relative wave velocity (or refractive index) $n=1$, with the front and back surface of the slab located 5 pixels away from the simulation boundaries in the $x$ direction. The slab itself is constructed by low-pass filtering a random refractive index distribution with values between 1 and 2 with a small imaginary part with a cut-off at 1/wavelength. The values are normalised between 1 and 2 after low-pass filtering. We used a truncation size of $t=8$ in the simulations unless stated otherwise.
We start by validating the domain decomposition framework on a small sample and then analyse the accuracy and convergence of the domain decomposition approach for different truncation sizes $t$ and different numbers of subdomains. Finally, we demonstrate domain decomposition in a large structure of $3.1\cdot 10^{7}$ cubic wavelengths.
@@ 4.1 Domain decomposition validation
Figure 5: Simulation results with and without domain decomposition for a 3D disordered medium of size $50\!\times\!50\!\times\!$ 50 × 50 wavelengths. 2D slices of (a) the random refractive index distribution used in the simulation, (b) the result for a single domain, and (c) the result with three subdomains in the $x$ direction, with the dashed grey lines indicating the subdomain boundaries. (d) Residual vs. iteration number for the simulations in b (solid line) and c (dashed line).
Fig. 5a shows a slice of the 3D refractive index distribution used in this test. The results of the simulations are shown in Fig. 5b and for the single domain and in Fig. 5c for the domain decomposed into three subdomains along the x-axis. The relative error between the two is just $1.5\cdot 10^{-4}$, confirming the validity of our domain decomposition approach.
Fig. 5d shows the residual as a function of the iteration number for the two simulations. The single-domain simulation converged in 500 iterations in 4.3 seconds, while the 3-domain simulation converged in 959 iterations in 6.1 seconds. The 1.9×1.9$\times$1.9 × increase in the number of iterations is due to a lower scaling coefficient $c$, with $a_{d}=2$ (Eq. (9)). Although this is a significant increase in the number of iterations, this factor is independent of the number of subdomains along an axis, as we will analyse in more detail in Section 4.2.2.
4.2 Convergence dependence on truncation parameter and number of subdomains simulation parameters
@@ 4.2.1 Dependence on truncation parameter
First, we examine the convergence behaviour as a function of $t$, the number of correction points used to generate the correction matrices $C_{11}$ and $A_{12}$. We decomposed the problem into two subdomains, varying $t$ from 00, indicating no correction, to $40$. We compare the relative error of these simulations with a single-domain simulation, as shown in Fig. 6a. The relative error between the two decays with more correction points, and with just $t=3$, the error is already less than $0.1\%$.
Figure 6: Convergence of a 3D simulation with two subdomains for different numbers of correction points $t$. (a) Relative error with a 1-domain simulation decays with $t$. (b) The number of iterations is almost constant for $t=3$ and higher, showing no dependence on $t$. (c) The total simulation time in seconds consistently increases for higher correction points.
Fig. 6b shows the number of iterations for convergence as a function of the number of correction points, plotted for $t=3$ and higher, as the accuracy is low for $t<3$. The figure shows that the number of iterations does not depend on $t$. The time for convergence (Fig. 6c) consistently increases with more correction points, which is expected since the size of the correction matrix increases.
@@ 4.2.2 Dependence on number of subdomains
Next, we examine the number of iterations and time as a function of the number of subdomains (Fig. 7). Fig 7a shows the number of iterations for different numbers of subdomains along the x and y-axis. The lowest number of iterations (728) is for a 1-domain simulation (dark blue pixel in the lower left corner in Fig. 7a). This case corresponds to a scaling coefficient $c$ based on only a wrapping correction along the $x$ axis, for which the system is non-periodic. When there is more than one subdomain along either the x or y-axis, the scaling factor $c$ reduces, which increases the iterations to $959$ with more domains along the x-axis (light blue band of pixels for $y=1$ in Fig. 7a), and $1068$ for more domains along the y-axis (light green band of pixels for $x=1$ in Fig. 7a). For more than one subdomain in both the x and y direction, scaling coefficient $c$ is even lower, increasing the iterations to around $1301$ (Fig. 7a, the dark red pixels for $x$ and $y>1$). Thus, the number of iterations increases only when subdomains are added along a new axis. The number of iterations is not dependent on the number of subdomains beyond one along an axis.
@@
Fig. 7b shows the simulation time in seconds as a function of the number of subdomains along the x and y-axis. For one domain in x or y directions, the simulation time is consistently low (dark blue band of pixels along the x and y axes at $x=1$ and $y=1$). It consistently increases with more subdomains due to an increase in the amount of communication between them. The time is highest for $10$ domains along both x and y directions, i.e., the total number of subdomains = $100$ (top right corner of Fig. 7b).

### v2

@@ Abstract
The modified Born series method is a fast and accurate method for simulating wave propagation in complex structures. Currently, its main limitation is that the simulation size is limited by the working memory of a single computer or graphics processing unit (GPU). Here, we present a domain decomposition method that removes this limitation. We show how to decompose large problems over subdomains while maintaining the method’s accuracy, memory efficiency, and guaranteed monotonic convergence. We demonstrate our approach by solving the Helmholtz problem for a complex structure of size $315\times 315\times 315$ wavelengths in just $1.4$ hours on a dual-GPU system.
@@
Helmholtz equation , Born series , modified Born series , Domain decomposition
@@ 1 Introduction
Osnabrugge et al. [3] showed that the MBS could be executed on a graphics processing unit (GPU) efficiently, solving Maxwell’s equations in a structure of $28\times 28\times 28$ wavelengths in just 3 seconds, representing a factor of $50$ speed-up compared to performing the simulations on a central processing unit (CPU). Recently, Valantinas and Vettenburg [26] solved the Helmholtz equation for a system of $2.1\cdot 10^{7}$ cubic wavelengths by mapping the steps of the MBS to layers in an artificial neural network framework, leveraging the extensive software and hardware acceleration available for these applications. However, GPUs have a limited amount of expensive working memory. They cannot be expanded as flexibly as the main working memory, adding a severe limitation to the simulation size [27, 28].
@@
Here, we present a domain decomposition approach for the MBS to solve the Helmholtz equation. Our approach, based on a non-overlapping domain decomposition strategy [33, 34, 35, 36], maintains the key benefits of MBS: low memory use, high accuracy, and guaranteed convergence, while minimising the overhead on memory use and computation time. Whereas it is common to explicitly formulate boundary conditions for the wave equation [36, 37, 38], our approach implicitly handles both the external boundary conditions and the boundary conditions at the domain boundaries, allowing both periodic and non-periodic boundaries. At the exterior, these boundaries are combined with a thin absorbing anti-reflection layer [3], eliminating the need for computationally expensive padding with thick absorbing boundary layers [8, 37].
We demonstrate the domain decomposition approach on a dual-GPU system through a large 3D simulation of $3.1\cdot 10^{7}$ cubic wavelengths, a size that is 1.93×1.93$\times$1.93 × larger than the maximum that could be accommodated on a single GPU and took only $1.4$ hours to complete. An open-source implementation in Python is available on GitHub [39].
@@ 2 Modified Born series
The key step in both the modified [2] and original [41] Born series formalisms is to split $A=L+V$, with $L\coloneqq c\left[\nabla^{2}+k_{0}^{2}\right]$ corresponding to wave propagation in a homogeneous medium with background wavenumber $k_{0}$, and $V=c\left[{k}^{2}(\boldsymbol{r})-k_{0}^{2}\right]$ the scattering potential.
@@
In contrast to the original Born series, which diverges for large or strongly scattering media [45], the MBS converges monotonically to the solution of Eq. (1) provided that $\alpha\leq 1$ and $ic$ is a real number less than $1/|k^{2}(\boldsymbol{r})-k_{0}^{2}|_{\infty}$ [2], where $|\cdot|_{\infty}$ denotes the maximum value over all $\boldsymbol{r}$. We can truncate this iteration when the norm of the residual $\Gamma^{-1}\left(\boldsymbol{y}-A\boldsymbol{x}^{(n)}\right)$ drops below a certain threshold.
@@
The splitting $A=L+V$ is visualised in Fig. 1a-c, where the operators are represented as matrices for illustrative purposes. In practice, these operators are not implemented as matrices since the full matrices would be too large to store in memory. The alternative of using sparse matrices would require replacing $L$ by a matrix with only a few diagonals, which is precisely the finite difference approximation that causes numerical dispersion and thus should be avoided. Instead, $L$ can be implemented as a convolution with the angular spectrum kernel [46, 47]
@@
where $\mathcal{F}$ and ${\mathcal{F}}^{-1}$ denote the forward and inverse Fourier transform, respectively, and $\boldsymbol{p}$ denotes the Fourier-space coordinate vector. In practice, a fast Fourier transform is used to evaluate this convolution efficiently. As shown in Fig. 1d, this approach does introduce wrapping artefacts. These artefacts can be significantly reduced by using an alternating sequence of offset Fourier transforms [3] and eliminated using our proposed method (see Section 3). Either of these methods is then combined with a thin absorbing boundary layer to ensure convergence [3].
@@
Figure 1: Matrix representation of splitting operator $A$ into $L+V$ for a 1D system. (a) Operator $A$ is split into (b) $L$ containing the Laplacian operator, a band along the diagonal, and (c) $V$, a diagonal matrix containing the scattering potential. (d) The operator $L$ with wraparound artefacts resulting from implementing $L$ as a fast convolution (Eq. (5)).
@@ 3 Domain decomposition of the modified Born series
An essential generalisation of the MBS introduced in [40] is that we have complete freedom to choose the splitting $A=L+V$, as long as the scaling factor $c$ can be selected such that $\lVert V\rVert<1$. We use this freedom to implement our domain decomposition framework.
This freedom enables us to implement $L$ as a fast convolution over each subdomain (see Fig. 2b). Representing $L$ and $V$ in block matrix form
$L=\begin{bmatrix}L_{11}&0\\ 0&L_{22}\end{bmatrix},\quad V=\begin{bmatrix}A_{11}-L_{11}&A_{12}\\ A_{21}&A_{22}-L_{22}\end{bmatrix},$ (8)
we see that $L$ is a domain-local operator and $V$ is mostly domain-local, with small, explicit contributions for transfer between domains (represented by $A_{12}$ and $A_{21}$, also seen in Fig. 2c). It is important to note that the system being solved here is still the same as in Eq. (2).
In $L$ (Eq. (8)), the diagonal blocks $L_{11}$ and $L_{22}$ (visualized in Fig. 2b) contain wraparound artefacts, similar to Fig. 1d. These wraparound artefacts appear in $V$ with the opposite sign, as $V\coloneqq A-L$. $V$ now has three contributions (Fig. 2c). The domain-local part constitutes the original scattering potential on the diagonal and the correction for the wrapping artefacts caused by the fast convolution, which we call $C_{11}$ and $C_{22}$. The non-local part of $V$ is the off-diagonal blocks $A_{12}$ and $A_{21}$ representing communication between the subdomains.
The interior subdomain communication and exterior boundary corrections are based on these matrices $C_{11}$, $C_{22}$, $A_{12}$, and $A_{21}$ combined with a thin absorbing boundary layer [3] on the exterior. For the exterior boundary, the matrices $C_{11}$ and $C_{22}$ are used to tackle the wrapping artefacts, which allows for only a thin absorbing boundary layer to minimize reflection artefacts. For the interior boundaries, we use the matrices $C_{11}$ and $C_{22}$ to tackle the wrapping artefacts at the interface between subdomains, and the matrices $A_{12}$ and $A_{21}$ for communication between the subdomains.
Thus, our approach does not require explicit steps to guarantee continuity at the subdomain boundaries but instead results in the naturally occurring boundary conditions at the interface through the matrices $A_{12}$ and $A_{21}$. The matrices $C_{11}$, $C_{22}$, $A_{12}$, and $A_{21}$ can all be constructed directly by comparing two situations: 1) the angular spectrum kernel applied to a point source using a linear convolution (Fig. 3a), and 2) the same kernel applied using a fast convolution on a subdomain (Fig. 3b). The difference between these two curves gives the required wrapping correction (which corresponds to a column of matrix $C_{11}$) and the transfer to the second subdomain (a column of matrix $A_{12}$). By repeating this process for different positions of the point source and thus different columns of the matrices, we recover matrices $C_{11}$ and $A_{12}$. Due to the symmetry of the angular spectrum kernel, the bottom-left part of matrix $C_{11}$ is equal to $-A_{12}$. The top-right part of matrix $C_{11}$ is given by the transpose $-A_{12}^{T}$.
@@
Figure 3c visualizes the bottom-left part of the domain-transfer matrix $A_{12}$ (also see Fig. 2c), which shows that the magnitude of the elements decreases rapidly away from the subdomain border. Therefore, we can truncate matrices $C$ and off-diagonal blocks of $A$ to only consider $t\ll N$ edge points near the boundaries. This way, the overhead of these additional computations is minimised.
Figure 4: Using truncated corrections in $V$ to reconstruct $A$. (a) Matrix representation of $A$ decomposed over two subdomains, demarcated by the dashed grey lines, but with truncated corrections. (b) $L$ decomposed over the subdomains remains unchanged and is the same as in Fig.2b. (c) $V$ with truncated corrections for wrapping and transfer, visible as square blocks of side $t=6$ in the corners of subdomain blocks. Note that the colour scale is logarithmic in both the positive and negative directions from 0.
@@ 3.1 Implementation
The domain decomposition framework is incorporated into the preconditioned Richardson iteration Eq. (6), but with the operators $L$ and $V$ computed for and applied to each subdomain. Thus, the fast convolution with ${(L+I)}^{-1}$ is computed for each subdomain $s$ as ${(L_{s}+I)}^{-1}$. It is the same for all subdomains of the same size and needs to be computed only once in that case. As discussed in Section 3, operator $V$ has three contributions, i.e. the scattering potential on the diagonal, which is now applied over the subdomains, the wrapping corrections $C$, and the transfer corrections (off-diagonal blocks of $A$). The steps for computing and applying the operator $V$ for every subdomain $s$ are as below:
@@
With this activation strategy, for example, a 3D simulation of size $100\times 100\times 100$ wavelengths split into $3\times 1\times 1$ subdomains was $12\%$ faster than without this strategy.
@@ 4 Results and Discussion
In this section, we demonstrate the domain decomposition of the MBS in 3D simulations. For all simulations in this section, we simulate light propagation through a scattering structure illuminated with a planar source located at $x=0$. We use absorbing boundaries with a thickness of 5 wavelengths [3], and a sampling of 4 points per wavelength.
The structure consists of tightly packed spheres, with a relative wave velocity (or refractive index) of 1.33 and a small imaginary part, distributed randomly in a medium with a refractive index $n=1$. The spheres were generated using the random_spheres() function in PoreSpy [48]. Unless stated otherwise, we used spheres of radius $3$ wavelengths and a truncation size of $t=8$ in the simulations.
We start by validating the domain decomposition framework on a small sample and then analyse the accuracy and convergence of the domain decomposition approach for different truncation sizes $t$ and numbers of subdomains. Finally, we demonstrate domain decomposition in a large structure of $3.1\cdot 10^{7}$ cubic wavelengths.
@@ 4.1 Domain decomposition validation
Figure 5: Simulation results with and without domain decomposition for a 3D medium of size $50\times 50\times 50$ wavelengths. 2D slices at z/2 of (a) the refractive index distribution used in the simulation, (b) the result for a single domain, and (c) the result with three subdomains in the x direction, with the dash-dotted black lines indicating the subdomain boundaries. (d) Residual vs. iteration number for the simulations in b (solid blue line) and c (dashed black line).
Fig. 5a shows a slice, through the centre of the z-axis, of the 3D refractive index distribution used in this test. The results of the simulations are shown in Fig. 5b for the single domain and in Fig. 5c for the domain decomposed into three subdomains along the x-axis. The relative error between the two is just $1.7\cdot 10^{-4}$, confirming the validity of our domain decomposition approach.
Fig. 5d shows the residual as a function of the iteration number for the two simulations. It can be seen that the residual decreases monotonically, which is a key property of the MBS. The single-domain simulation converged in 409 iterations in 2.3 seconds, while the 3-domain simulation converged in 2,908 iterations in 11.7 seconds. The 7×7$\times$7 × increase in the number of iterations is due to a lower scaling coefficient $c$, with $a_{d}=2$ (Eq. (9)). Although this is a significant increase in the number of iterations, this factor is independent of the number of subdomains along an axis, as we will analyse in more detail in Section 4.2.2.
4.2 Convergence dependence on truncation parameter and number of subdomains
@@ 4.2.1 Dependence on truncation parameter
First, we examine the convergence behaviour as a function of $t$, the number of correction points used to generate the correction matrices $C_{11}$ and $A_{12}$ (Section 3). We decomposed the problem in Section 4.1 into two subdomains along the x-axis and varied $t$ from 00, indicating no correction, to $40$. We compare the relative error of these simulations with a single-domain simulation, as shown in Fig. 6a. The relative error between the two decays with more correction points, and with just $t=4$, the error is already less than $0.1\%$.
Figure 6: Convergence of a 3D simulation with two subdomains for different numbers of correction points $t$. (a) Relative error with a 1-domain simulation decays with $t$. (b) The number of iterations is almost constant for $t=4$ and higher, showing no dependence on $t$. (c) The total simulation time in seconds consistently increases for more correction points.
Fig. 6b shows the number of iterations for convergence as a function of the number of correction points, plotted for $t=4$ and higher, as the accuracy is low for $t<4$. The figure shows that the number of iterations does not depend on $t$. The time for convergence (Fig. 6c) consistently increases with more correction points, which is expected since the size of the correction matrix increases.
@@ 4.2.2 Dependence on number of subdomains
Next, we examine the number of iterations and time as a function of the number of subdomains (Fig. 7). Fig 7a shows the number of iterations for different numbers of subdomains along the x and y-axis. The lowest number of iterations (1160) is for a 1-domain simulation (black pixel in the lower left corner in Fig. 7a). This case corresponds to a scaling coefficient $c$ based on only a wrapping correction along the x-axis, for which the system is non-periodic. When there is more than one subdomain along either the x or y-axis, the scaling factor $c$ reduces, which increases the iterations to around $1751$ with more domains along the x-axis (dark pink band of pixels for $y=1$ in Fig. 7a), and around $2018$ for more domains along the y-axis (orange band of pixels for $x=1$ in Fig. 7a). For more than one subdomain in both the x and y direction, scaling coefficient $c$ is even lower, increasing the iterations to around $2448$ (Fig. 7a, the light yellow pixels for $x$ and $y>1$). Thus, the number of iterations increases only when subdomains are added along a new axis. The number of iterations is not dependent on the number of subdomains beyond one along an axis.
@@
Fig. 7b shows the simulation time in seconds as a function of the number of subdomains along the x and y-axis. The simulation time is consistently low for one domain in x or y directions (black band of pixels along the x and y axes at $x=1$ or $y=1$). It consistently increases with more subdomains due to an increase in the amount of communication between them. The time is highest for $10$ domains along both x and y directions, i.e., the total number of subdomains = $100$ (top right corner of Fig. 7b).